### v1

# SPHERE: Spherical partitioning for large-scale routing optimization

Abstract
We study shortest-path routing in large weighted, undirected graphs, where expanding search frontiers raise time and memory costs for exact solvers. We propose SPHERE, a source-target-aware heuristic that identifies an $s$-$t$ overlap: vertices that are close to both $s$ and $t$ in hop count. Selecting an anchor $a$ in this overlap partitions the task into two subproblems with unchanged problem-topology , $s\to a$ and $a\to t$; if either remains large, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating the resulting subpaths yields a valid $s\to t$ route without boundary repair. SPHERE is independent of the downstream solver (e.g., Dijkstra) and exposes parallelism across subproblems. On large networks, it achieves faster runtimes and smaller optimality gaps than Louvain-based routing and a METIS-based pipeline, even on graphs with more than a million nodes and edges, while also outperforming Dijkstra in runtime.
1 Introduction
Point-to-point routing in large weighted, undirected graphs is a canonical problem in operations research, one that emerged alongside the field’s inception.. Exact label-setting or label-correcting methods (e.g., Dijkstra) can certify optimality (Dijkstra, 1959), but their search frontiers grow quickly on large networks and become memory- and time-intensive. Graph partitioning methods (e.g., Kernighan-Lin, multilevel partitioning/METIS, spectral cuts) reduce problem size by cutting the graph into parts (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000), yet cuts chosen without regard to a specific query $(s,t)$ can intersect narrow passages that many feasible sstt routes must use, which can exclude high-quality routes or require costly boundary repair. Bidirectional search reduces depth by growing frontiers from $s$ and $t$ (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016), but still explores a single coupled state space and may expand many partial routes on large graphs. For completeness, we note that while feasibility for unconstrained shortest paths is trivial on connected graphs, constrained variants (e.g., resource- or budget-limited paths) are NP-hard; our focus here is the unconstrained case with nonnegative weights, which admits an optimal solution via Dijkstra’s algorithm (Dijkstra, 1959).
We propose SPHERE, a query-aware partition that aligns computation to the sstt geometry. Let $B_{s}(r)$ and $B_{t}(r)$ denote hop-distance spheres around $s$ and $t$. SPHERE starts from a radius tuple $(r_{s},r_{t})$ with guaranteed overlap $B_{s}(r_{s})\cap B_{t}(r_{t})\neq\emptyset$, then applies a monotone decrement rule that reduces the radii (when needed, both simultanously) so that $|r_{s}-r_{t}|\leq 1$ while preserving overlap, and selects the minimal balanced pair, i.e. no further admissible unit decrement of either radius (still keeping $|r_{s}-r_{t}|\leq 1$) leaves the intersection nonempty. Then it selects an anchor $a\in B_{s}(r_{s})\cap B_{t}(r_{t})$ and splits the query into $s\!\to\!a$ and $a\!\to\!t$, solved on the corresponding induced subgraphs under a tunable size cap. If a subproblem still exceeds the cap, the procedure recurses on its induced subgraph. Because the cut lies inside the overlap, concatenating subpaths at $a$ yields a valid sstt route without boundary repair. The partition is solver-agnostic and the independent subproblems can be solved in parallel.
Our contributions are:
  1. 1.
    Query-aware partition. We introduce a simple heuristic that adaptively splits the graph based on the geometry of the source-target pair, ensuring that each part preserves connectivity and can be rejoined without additional repair.
  2. 2.
    Parallel subinstances. The subproblems generated by SPHERE are independent induced subgraphs, naturally exposing parallelism and enabling scalable routing on distributed systems with effort scaling to hardware.
  3. 3.
    Empirical validation. On large road graphs, SPHERE achieves faster runtimes than Louvain-based routing and a METIS-based pipeline while matching or improving route quality, benefiting further from inherent parallelism.
SPHERE is complementary to downstream solvers and can wrap Dijkstra or A* (Dijkstra, 1959; Hart et al., 1968), as well as modern speedup indices such as ALT/landmarks (Goldberg and Harrelson, 2005), reach-based pruning (Gutman, 2004), contraction hierarchies (Geisberger et al., 2008), hub labeling (Abraham et al., 2011), and transit-node routing (Bast et al., 2016); we evaluate with common solvers and baselines and discuss integration in Section 5. Next, we motivate these design choices by examining where existing approaches trade-off feasibility, and instance size, before positioning SPHERE within the literature and formalizing our approach.
2 Motivation
Routing at scale must meet three criteria simultaneously: control per-instance complexity, make an $s$-$t$ route feasible, and remain compatible with trusted downstream solvers while enabling parallel execution. Existing tools typically sacrifice at least one of these: Labeling methods (Feillet et al., 2004; Irnich and Desaulniers, 2005) preserve feasibility but generate a combinatorial explosion of partial routes on large graphs. Graph partitioners (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Shi and Malik, 2000) control instance size, but because their boundaries are chosen without regard to $s$ and $t$, they may cut across a bottleneck that every feasible sstt path must traverse, risking the loss of valid routes. Bidirectional search (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997; Holte et al., 2016) reduces search depth by expanding from both ends, $s$ and $t$, but the two frontiers still form a single combined search space and therefore cannot be partitioned into smaller, independent pieces suitable for parallel solution.
SPHERE overcomes these trade-offs by aligning the partition with the specific request. To ensure controllable size, it expands hop-distance spheres from $s$ and $t$ until reaching the last nonempty overlap, where a user-defined radius limits the diameter of each subproblem. Second, for feasibility, SPHERE selects an anchor $a$ in the overlap and solves $s\!\to\!a$ and $a\!\to\!t$ on their respective induced subgraphs, so cuts always lie inside the overlap and concatenation at $a$ is guaranteed to generate a feasible sts-t route. This preserves all local connectivity within each part and ensures that every feasible subpath remains feasible in the global graph. Third, the resulting subproblems are independent, enabling parallel execution across multiple cores, and the design is solver-agnostic which means that any downstream optimizer, from exact Mixed-Integer Programming (MIP) on small pieces to strong metaheuristics on larger ones (Pisinger and Ropke, 2007; Vidal, 2022), can be applied without modification. In short, SPHERE offers a partition strategy that controls complexity, preserves feasibility, and exposes natural parallelism while remaining compatible with the solvers practitioners already trust.
3 Related Work
Point-to-point routing on large weighted graphs has been studied from several perspectives. Classical exact methods such as label-setting and label-correcting algorithms (e.g., Dijkstra) can guarantee global optimality on connected graphs with nonnegative weights (Dijkstra, 1959). Dynamic-programming and labeling variants extend this to richer settings and can certify optimality, but they often face an explosion in the number of labels on large graphs (Feillet et al., 2004; Irnich and Desaulniers, 2005). Approximation schemes exist for restricted variants (Hassin, 1992), trading accuracy for speed, yet these techniques operate on the full graph and therefore do not reduce the instance size before optimization.
Graph partitioning offers a complementary idea: simplify the problem by splitting the graph. Kernighan-Lin improves an initial cut by greedy node swaps to reduce the number of edges crossing between parts while keeping parts balanced (Kernighan and Lin, 1970). Multilevel schemes such as METIS repeatedly coarsen the graph, compute a cut on the coarse instance, and refine it during uncoarsening, which gives strong scalability in practice (Karypis and Kumar, 1998). In routing pipelines, these partitions are often used to build a quotient graph for a coarse path that is later refined within parts. A limitation of all such methods is that the cut is chosen without regard to the specific start-target pair $(s,t)$, so the boundary can lie across a narrow passage that every feasible $s$-$t$ path must use. Once this happens, the coarse or restricted routing stage may miss or degrade the necessary connection.
Bidirectional search reduces search depth by expanding frontiers from both ends and meeting near the middle (Nicholson, 1966; Pohl, 1969; Kaindl and Kainz, 1997). The MM algorithm of Holte et al. (2016) provides a priority rule that ensures expansions do not go beyond the midpoint of the optimal solution. These methods improve exploration efficiency but still operate within a single coupled search space and do not produce small, independent instances. In large graphs this can still generate many partial routes, and unlike partition-based methods, they do not offer a straightforward path to parallel, solver-agnostic execution.
4 Spherical Partitioning
4.1 Spherical subgraphs
We begin by fixing distance and the local “window” around a node. In the following, let every graph $G=(V,E)$ be a finite, connected, undirected graph with positive edge weights. Unless stated otherwise we measure distance in hops (unweighted shortest path); all statements below remain valid for the weighted shortest-path distance.
Definition 1 (Spheres).
For $R\in\mathbb{N}$ and a reference node $\tilde{v}\in V$, the (closed) sphere is
$S_{R}(\tilde{v})\;:=\;\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}\subseteq V,$
where $d_{G}(\tilde{v},v)$ is the unweighted shortest path distance between $\tilde{v}$ and $v$.
[…]
Definition 2 (Spherical subgraph).
[…]
$E^{\mathrm{sp}}_{R}(\tilde{v})=\bigcup_{v\in S_{R}(\tilde{v})}\mathcal{E}\!\bigl(\mathrm{SP}(\tilde{v},v)\bigr),$
[…]
Definition 3 (Induced spherical subgraph).
[…]
4.2 Partition cut
With these building blocks in place, we now align the local windows to a specific query $(s,t)$. For radii $(R_{s},R_{t})\in\mathbb{N}^{2}$, define the overlap
$O(R_{s},R_{t})\;:=\;S_{R_{s}}(s)\cap S_{R_{t}}(t).$
We control how radii evolve, how they are initialized, and how an anchor is chosen in the overlap.
Definition 4 (Decrement rule).
A decrement rule is a map $\mathrm{decru}:\mathbb{N}^{2}\to\mathbb{N}^{2}$ that is monotone in the lexicographic order, i.e.,
$(x_{1},x_{2})\leq_{\mathrm{lex}}(y_{1},y_{2})\;\Longleftrightarrow$ $\begin{aligned} &x_{1}<y_{1},\\ &\text{or }(x_{1}=y_{1}\text{ and }x_{2}\leq y_{2}).\end{aligned}$
and never increases any coordinate.
Example (monotone decrement).
A simple rule that satisfies the definition is
$\mathrm{decru}(R_{s},R_{t})=\begin{cases}(R_{s}{-}1,R_{t}),&R_{s}\geq R_{t},\\[2.0pt] (R_{s},R_{t}{-}1),&\text{otherwise.}\end{cases}$
It is coordinatewise nonincreasing, and it is lex-monotone: if $(R_{s},R_{t})\leq_{\mathrm{lex}}(R^{\prime}_{s},R^{\prime}_{t})$, then either $R_{s}<R^{\prime}_{s}$ (so the first branch cannot produce a larger first coordinate) or $R_{s}=R^{\prime}_{s}$ with $R_{t}\leq R^{\prime}_{t}$, in which case both images either decrement the same coordinate or the smaller pair cannot “overtake” the larger one in lexicographic order.
The map $\mathrm{decru}$ from Definition 4 determines how radii shrink once chosen. To make the procedure operative, we also need: (i) an initial choice of radii that guarantees the spheres overlap at the start, and (ii) a way to convert any nonempty overlap into a concrete split point. These are captured by the following two rules.
Definition 5 (Starting rule).
A starting rule $\mathrm{staru}:V\times V\to\mathbb{N}^{2}$ returns $(R_{s}^{0},R_{t}^{0})$ with nonempty overlap $O(R_{s}^{0},R_{t}^{0})\neq\emptyset$ (e.g., $R_{s}^{0}{+}R_{t}^{0}\geq d_{G}(s,t)$; a stronger but safe choice is $S_{R_{s}^{0}}(s)\cup S_{R_{t}^{0}}(t)=V$).
Definition 6 (Anchor rule).
An anchor rule $\mathrm{anru}:\mathcal{P}(V)\to V$ selects $a\in O(R_{s},R_{t})$ (e.g., uniform random).
Bringing these pieces together, the three rules: starting rule for initialization (Definition 5), decrement rule for shrinking (Definition 4), and anchor rule for selecting the split point (Definition 6), jointly define the following algorithm, which we refer to as the Partition cut:
Algorithm 1 Partition cut
1:graph $G=(V,E)$, terminals $(s,t)$,
2:     rules $(\mathrm{staru},\mathrm{decru},\mathrm{anru})$
3:  radii $(\bar{R}_{s},\bar{R}_{t})$, overlap $O$, anchor $a$
[…]
5:while $O(R_{s},R_{t})\neq\emptyset$ do
6:  $(\bar{R}_{s},\bar{R}_{t})\leftarrow(R_{s},R_{t})$
7:  $(R_{s},R_{t})\leftarrow\mathrm{decru}(R_{s},R_{t})$
8:end while
[…]
10:return $(\bar{R}_{s},\bar{R}_{t},O,a)$
The intuition behind Algorithm 1 can be explained as follows and is visualized in Figure 2. Grow spheres around $s$ and $t$ until they overlap, then shrink via $\mathrm{decru}$ until the next shrink would empty the overlap. Pick an anchor $a$ according to Definition 6 in this last nonempty overlap, and split the query into $s\!\to\!a$ and $a\!\to\!t$ subproblems on the corresponding induced spheres.
[…]
Figure 2: Last nonempty overlap. Blue and red spheres overlap in the green region containing both $a$ and $a^{\prime}$. Potential anchor node $a$ is selected according to Definition 6.
A natural question is whether the Partition cut of Algorithm 1 actually preserves feasibility of the original $s{\to}t$ routing problem. In other words, does splicing the two subpaths at the chosen anchor still yield a valid global route? The following statement makes this precise:
Proposition 4.1 (Feasibility preservation).
Let $(\bar{R}_{s},\bar{R}_{t},O,a)$ be returned by Algorithm 1. If $P_{s\to a}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{s}}(s)$ and $P_{a\to t}$ is a feasible path in $G^{\mathrm{ind}}_{\bar{R}_{t}}(t)$, then $P_{s\to a}\!\circ\!P_{a\to t}$ is a feasible $s\to t$ path in $G$.
Proof sketch.
Both subgraphs are induced, hence every used edge is in $G$. Since $a\in O(\bar{R}_{s},\bar{R}_{t})$, concatenating at $a$ yields a valid $s\to t$ path in $G$. ∎
4.3 Recursive spherical partitioning
The following algorithm extends the partition cut from Algorithm 1 into a recursive partitioning routine. Its goal is to partition a global $s\!\to\!t$ routing task into a set of smaller subproblems, each confined to an induced sphere whose radius does not exceed a user-defined bound $R_{\max}$. These subproblems are returned as task triples $(H,u,w)$, meaning: “route from $u$ to $w$ within the induced subgraph HH.”
Algorithm 2 Recursive spherical partitioning
1graph $G=(V,E)$, terminals $(s,t)$,
2     radius cap $R_{\max}$, rules $\mathrm{staru},\mathrm{decru},\mathrm{anru}$
3  set $\mathcal{T}$ of task triples $(H,u,w)$
4function SphPartition($G,s,t,R_{\max}$)
5  $\mathcal{T}\leftarrow\emptyset$, $\texttt{rule}\leftarrow(\mathrm{staru},\mathrm{decru},\mathrm{anru})$
6  $(R_{s},R_{t},O,a)\leftarrow\textsc{PartCut}(G,s,t,\texttt{rule})$ $\triangleright$ Alg. 1
7  if $R_{s}\leq R_{\max}$ then
8  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{s}}(s),\,s,\,a)\}$
9  else
10  $\mathcal{T}\leftarrow\mathcal{T}\cup\textsc{SphPartition}(G^{\mathrm{ind}}_{R_{s}}(s),\,s,\,a,\,R_{\max})$
11  end if
12  if $R_{t}\leq R_{\max}$ then
13  $\mathcal{T}\leftarrow\mathcal{T}\cup\{(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t)\}$
14  else
15  $\mathcal{T}\leftarrow\mathcal{T}\cup\textsc{SphPartition}(G^{\mathrm{ind}}_{R_{t}}(t),\,a,\,t,\,R_{\max})$
16  end if
17  return $\mathcal{T}$
18end function
Algorithm 2 starts by initializing the set of tasks $\mathcal{T}$ to be empty (lines 1-3). It then performs a partition cut on the original problem $(G,s,t)$, which returns the radii $(R_{s},R_{t})$, the last nonempty overlap $O$, and an anchor $a\in O$. The anchor acts as a splitting point that separates the global $s\!\to\!t$ route into two sides.
The algorithm then branches into two recursive checks. On the $s$-side (lines 4-7), the relevant subproblem is $(G^{\mathrm{ind}}_{R_{s}}(s),s,a)$. If the radius $R_{s}$ is already below the cap $R_{\max}$, this subproblem is considered sufficiently small and is added to the task set $\mathcal{T}$. If not, the procedure calls itself recursively on this induced subgraph to split it further. Symmetrically, the $t$-side branch (lines 8-11) processes the subproblem $(G^{\mathrm{ind}}_{R_{t}}(t),a,t)$: either it is recorded directly if $R_{t}\leq R_{\max}$, or it is further partitioned by another recursive call.
Finally, once both branches have been processed, the algorithm returns (line 12) the accumulated set $\mathcal{T}$ of task triples. By construction, concatenating the solutions to these recorded subproblems along their anchors yields a valid overall $s\!\to\!t$ route.
(G,s,t)(G,s,t)Partition cut $\Rightarrow$$(R_{s},R_{t},O,a)$$G^{\mathrm{ind}}_{R_{s}}(s),\ s\to a$$G^{\mathrm{ind}}_{R_{t}}(t),\ a\to t$if $R_{s}\leq R_{\max}$: record; else: recurseif $R_{t}\leq R_{\max}$: record; else: recurse
Figure 3: Recursive splitting. Any side above the radius cap is split again by Algorithm 1; leaves are independent subproblems.
Figure 3 illustrates this recursive structure. Each partition cut splits the problem into two boxes, one for the start-side and one for the target-side. If the radius on a side is within the threshold, the box is recorded as a leaf; otherwise, it is expanded by another partition cut.
4.3.1 Baseline application
The recursive partition in Algorithm 2 produces a set of leaf tasks, each represented by a triple $(H,u,w)$ where $H$ is an induced spherical subgraph and $u,w$ are its boundary terminals. At this stage, the global $s\to t$ routing problem has been reduced to a collection of independent local subproblems, each small enough to be solved efficiently by standard methods.
Algorithm 3 Baseline application with spherical partitioning
1:graph $G=(V,E)$, terminals $(s,t)$,
2:     radius cap $R_{\max}$, solver $\mathrm{optimize}(\cdot)$
3:  feasible route $\mathcal{R}$
[…]
5:for all $(H,u,w)\in\mathcal{T}$ do
6:  $\mathcal{R}\leftarrow\mathcal{R}\circ\mathrm{optimize}(H,u,w)$
7:end for
8:return $\mathcal{R}$
Algorithm 3 describes how to turn the recursive spherical partitioning into a full routing solution. First, the input graph $G$ with terminals $s,t$ is partitioned into a collection of smaller subproblems (line 3). Each subproblem $(H,u,w)$ is an induced subgraph $H$ together with its local terminals $u$ and $w$. Then, for each subproblem, an arbitrary downstream solver $\mathrm{optimize}(H,u,w)$ is applied to compute a local path (lines 4-5). These local solutions are concatenated in the order determined by the anchors of the partition, resulting in a valid global route $\mathcal{R}$ from $s$ to $t$ (line 6). Since the subproblems are independent, they can also be solved in parallel, and the procedure remains agnostic to the choice of solver.
5 Numerical Experiments
Our experiments are designed to evaluate whether our method delivers significant performance gains (in terms of runtime and optimality gap) over established graph-routing baselines.
5.1 Dataset
For our empirical evaluation we use the well-known West-USA-road benchmark suite from the 9th DIMACS Implementation Challenge on Shortest Paths (DIM, 2006). These graphs encode real-world road networks at the level of individual intersections and road segments. Each node corresponds to a road intersection or endpoint, and each edge represents a street segment with a positive weight equal to its physical length. The dataset is widely used in the routing literature as a benchmark for large-scale shortest-path computations, since it combines realistic geometry, heterogeneous edge weights, and large graph sizes (millions of nodes and edges).
5.2 Experimental setup
For each problem seed $p\in\{1,\dots,30\}$ we draw a source-target pair $(s,t)$ uniformly at random from the node set (rejecting $s=t$). For each $(s,t)$, we run five independent trials with inner seeds $q\in\{1,\dots,5\}$ to capture variability from randomized components. For instance, the anchors are selected randomly according to Definition 6. We report results per instance and aggregate by averaging across the $5$ inner seeds for each problem seed.
Given $(s,t)$, we apply Algorithm 2 and selects an anchor node uniformly at random from this set of nodes (the randomness is controlled by the inner seed $q$). The graph is partitioned into subgraphs around $s$ and $t$ until a maximum radius threshold is met. $R_{max}$ from Algortihm 3 is set to $1800$. Throughout development we found a robust rule of thumb: when routing within a spherical subgraph of hop radius $R$, choosing $R_{\max}\in[\,R/2,\;R/1.7\,]$ delivers high-quality solutions at both low computational cost and effort. For the West-USA instances, a single sphere with $R\approx 3200$ covers the region; thus $R_{\max}=1800$ lies squarely in this band, balancing overlap likelihood and leaf size. Each subgraph shortest-path subproblem is solved with Dijkstra’s algorithm. Because our spherical partitioning produces independent induced subgraphs, we parallelize these subgraph solves and report end-to-end wall-clock time (including partition cut and final concatenation). By contrast, the baselines do not naturally partition into independent leaf tasks so we do not apply the same solving subproblem parallelization to them.
Baselines.
As baselines, we first include an optimal reference obtained by running Dijkstra’s algorithm directly on the full graph. To compare against partition-based approaches, we also evaluate a corridor-routing strategy in which the graph is partitioned into $k=64$ cells using PyMetis; a path is computed on the resulting quotient graph, and Dijkstra’s algorithm is then applied within the induced corridor to recover a detailed route (Karypis and Kumar, 1998). In addition, we test a partition heuristic based on community detection via the Louvain method: routing is first performed on the community graph, and the solution is subsequently refined by applying Dijkstra on the induced subgraph spanned by the relevant communities (default parameters; seeded where applicable) (Blondel et al., 2008). The corridor-routing and Louvain pipelines introduce partitions, yet can not be solved independently and not conditioned on the specific $(s,t)$, so they likewise do not yield independent leaves that one could solve with a plug-in solver and then concatenate repair-free (e.g. fixing feasibility) (Karypis and Kumar, 1998; Blondel et al., 2008).
Metrics. Let $m\in\{\textsc{Sphere},\ \textsc{Corridor},\ \textsc{Louvain}\}$ denote the method. For each problem seed $p$ and inner seed $q$, let $C_{p,q}^{(m)}$ be the route cost returned by $m$, and let $C_{p}^{\star}$ be the shortest-path cost from Dijkstra’s algorithm on the full graph for the same $(s,t)$. Define the relative optimality gap
$\Delta_{p,q}^{(m)}\;=\;\frac{C_{p,q}^{(m)}-C_{p}^{\star}}{C_{p}^{\star}}.$
For each $p$, aggregate across the five inner seeds by the median:
$\widetilde{\Delta}_{p}^{(m)}\;=\;\operatorname{median}_{q=1,\dots,5}\,\Delta_{p,q}^{(m)},$
$\widetilde{T}_{p}^{(m)}\;=\;\operatorname{median}_{q=1,\dots,5}\,T_{p,q}^{(m)},$
where $T_{p,q}^{(m)}$ is the end-to-end wall-clock time. For completeness, we also report means:
$\overline{\Delta}_{p}^{(m)}\;=\;\tfrac{1}{5}\sum_{q=1}^{5}\Delta_{p,q}^{(m)},\qquad\overline{T}_{p}^{(m)}\;=\;\tfrac{1}{5}\sum_{q=1}^{5}T_{p,q}^{(m)}.$
5.3 Results
Across 30 West-USA instances with five inner seeds each, SPHERE achieves faster end-to-end runtimes than the three baselines and matches near-optimal route quality. We use performance and accuracy profiles for visualization (Dolan and Moré, 2002; Beiranvand et al., 2017). Both are computed per instance from the median over seeds. Mean-based profiles for runtime and accuracy are presented as well.
(a) Runtime profile based on the median over five seeds
(b) Accuracy profile based on the median over five seeds
(c) Runtime profile based on the mean over five seeds
(d) Accuracy profile based on the mean over five seeds
Figure 4: Performance and accuracy profiles of 30 instances based on median and mean over five seeds.
Runtime.
Figures 4(a) and 4(c) plot the fraction of instances solved within a performance ratio $\tau$ of the best method on that instance. SPHERE’s curve lies above the Louvain and corridor routes for most $\tau$, indicating uniformly stronger or equal runtime across instances. Interestingly, for small $\tau$, Louvain’s curve lies below Dijkstra’s, despite being a suboptimal method. However, at larger $\tau$ it crosses and stays above, indicating tail slowdowns for Dijkstra. This suggests that SPHERE not only improves median runtime, but also avoids the heavy-tailed slowdowns observed in both Dijkstra and Louvain, resulting in tighter runtime distributions and fewer extreme outliers, which makes SPHERE’s runtimes more predictable across different instances.
Route quality.
Figures 4(b) and 4(d) reports the fraction of instances whose relative optimality gap is at most $\tau$. Dijkstra attains zero gap by construction. SPHERE’s curve dominates the partition baselines across $\tau$, yielding smaller or equal gaps on the majority of instances. This confirms that cutting inside the $s$-$t$ overlap and solving on induced spheres preserves high-quality routes without boundary repair.
Pareto dominance
Against the partition baselines (Louvain and METIS corridor), SPHERE Pareto-dominates on 26 of 30 instances in terms of average metrics. It is both faster and has a smaller optimality gap (Appendix A.3). The remaining three show a speed-quality trade-off where SPHERE is not dominated by another algorithm. In other words, SPHERE consistently lies on or close to the empirical speed–quality frontier, while Louvain and corridor-routing often face a trade-off between being faster but less accurate, or more accurate but substantially slower.
Robustness and aggregation
Repeating the analysis with means over seeds yields profiles that do not differ materially from the median-based plots, supporting robustness to the randomized anchor selection. Additionally, the full results in the appendix show low dispersion across seeds. For SPHERE, the optimality-gap standard deviation is 0 on 16 instances and is the lowest among all methods on all but one instance. For further Information, we refer to Appendix A.3.
6 Conclusion and Outlook
We introduced SPHERE, a routing scheme that splits an sstt query by cutting inside the last nonempty overlap of hop-spheres around $s$ and $t$. This placement preserves feasibility without boundary repair, confines each side to an induced subgraph, and allows the two subproblems to be solved independently with standard shortest-path routines and in parallel.
On West-USA-road graphs, SPHERE achieves faster end-to-end times and equal or smaller optimality gaps than a Louvain-based route and a METIS-corridor pipeline on the large majority of instances; runtimes are also more predictable, as heavy-tail slowdowns seen for the baselines largely disappear. In contrast to bidirectional search, which still explores a single coupled state space, SPHERE yields separable leaves. In contrast to static partitions whose cuts ignore $(s,t)$ and can sever bottlenecks, cutting within the overlap keeps the route intact and avoids corridor repair.
Looking ahead, the main levers are the starting rule, the decrement rule, and the anchor selection. Even simple choices already perform well, but data-driven or adaptive variants should improve robustness across graphs and query distributions. A second priority is to tune the radius cap $R_{\max}$ automatically (per instance or region) to balance overlap likelihood against subproblem size. These steps retain the current simplicity and parallelism while targeting further speed and stability.
Appendix A Appendix
A.1 Notation
{longtblr}
[ caption = Notation, label = tab:notation ]colspec = @ Q[l,wd=4.6cm] X @, rowhead = 2 Symbol Description
$V=\{1,\dots,m\}$ Node (vertex) index set.
$\mathcal{P}(V)$ Power set of $V$.
$E\subseteq V\times V$ Edge set; undirected unless stated.
$G=(V,E)$ Graph with node set $V$ and edge set $E$.
$d_{G}(u,v)$ Shortest unweighted path distance on $G$.
$G[U]$ Induced subgraph on node set $U$.
$S_{R}(\tilde{v})=\{\,v\in V:\ d_{G}(\tilde{v},v)\leq R\,\}$ (Closed) hop sphere (ball) of radius $R$ around $\tilde{v}$.
$S_{s}(r)\equiv S_{r}(s),\;S_{t}(r)\equiv S_{r}(t)$ Shorthand spheres around $s$ and $t$.
$O(R_{s},R_{t})=S_{R_{s}}(s)\cap S_{R_{t}}(t)$ Overlap of the two spheres.
$a\in V$,  $S^{\dagger}\subseteq O(R_{s},R_{t})$ Anchor node; optional candidate anchor set.
$G_{s}=G[S_{R_{s}}(s)],\;G_{t}=G[S_{R_{t}}(t)]$ Induced subgraphs around $s$ and $t$ at radii $R_{s},R_{t}$.
$(H,u,w)$,  $\mathcal{T}$ Task triple “route $u\to w$ inside HH”; set of leaf tasks.
$|X|$ Cardinality of a (finite) set $X$.
A.2 Parameters
{longtblr}
[ caption = Parameters, label = tab:params ]colspec = @ Q[l,wd=3.4cm] X @, rowhead = 2 Symbol Description
$R_{\max}\in\mathbb{N}$ Maximum allowed radius for spherical subgraphs (controls subproblem diameter).
$V_{\max}\in\mathbb{N}$,  $E_{\max}\in\mathbb{N}$ Optional caps on nodes/edges permitted in any subgraph $G[U]$.
$\varepsilon_{\text{overlap}}\in\mathbb{N}_{0}$ Minimum required overlap size (require $|O(R_{s},R_{t})|\geq\varepsilon_{\text{overlap}}$).
$\Delta r\in\mathbb{N}$ Radius increment/decrement step during sphere growth/shrinking (if using fixed steps; typically $\Delta r=1$).
$L_{\max}\in\mathbb{N}$ Maximum recursion depth for repeated splits.
$k_{\text{anchor}}\in\mathbb{N}$ Number of anchors to select from $O$ (often $k_{\text{anchor}}=1$).
A.3 Performance per instance
Table 1: WestUSA – Routing
avg compute time avg opt. gap median opt. gap std opt. gap
P1 SPHERE 135.73 0.08 0.06 0.02
Corridor Route 173.97 0.24 0.07 0.36
Louvain 1047.77 0.25 0.28 0.05
P2 SPHERE 138.27 0.12 0.12 0.00
Corridor Route 187.87 0.14 0.09 0.16
Louvain 1072.17 0.34 0.35 0.05
P3 SPHERE 165.25 0.03 0.03 0.00
Corridor Route 190.12 0.08 0.10 0.06
Louvain 1130.18 0.18 0.18 0.05
P4 SPHERE 370.06 0.11 0.11 0.00
Corridor Route 574.18 0.25 0.26 0.10
Louvain 1230.71 0.37 0.37 0.03
P5 SPHERE 126.88 0.05 0.05 0.00
Corridor Route 171.35 0.10 0.11 0.06
Louvain 1129.89 0.14 0.14 0.06
P6 SPHERE 301.07 0.02 0.03 0.00
Corridor Route 571.56 0.18 0.15 0.14
Louvain 1265.49 0.25 0.20 0.12
P7 SPHERE 213.59 0.00 0.00 0.00
Corridor Route 354.04 0.27 0.36 0.17
Louvain 1197.91 0.32 0.32 0.04
P8 SPHERE 157.02 0.01 0.01 0.00
Corridor Route 224.89 0.32 0.27 0.16
Louvain 1165.53 0.09 0.07 0.03
P9 SPHERE 390.68 0.02 0.02 0.00
Corridor Route 396.99 0.13 0.10 0.10
Louvain 1274.68 0.23 0.21 0.06
P10 SPHERE 137.20 0.02 0.02 0.00
Corridor Route 197.31 0.02 0.00 0.03
Louvain 1141.72 0.27 0.33 0.10
P11 SPHERE 130.96 0.13 0.13 0.01
Corridor Route 167.76 0.20 0.07 0.24
Louvain 1057.36 0.33 0.33 0.00
P12 SPHERE 264.38 0.04 0.04 0.00
Corridor Route 234.53 0.24 0.20 0.13
Louvain 1160.17 0.25 0.17 0.13
P13 SPHERE 123.22 0.00 0.00 0.00
Corridor Route 200.92 0.10 0.00 0.16
Louvain 1308.92 0.18 0.19 0.03
P14 SPHERE 234.34 0.14 0.14 0.00
Corridor Route 358.43 0.25 0.24 0.17
Louvain 1414.78 0.33 0.31 0.05
P15 SPHERE 211.37 0.00 0.00 0.00
Corridor Route 300.88 0.11 0.10 0.09
Louvain 1288.33 0.15 0.15 0.02
P16 SPHERE 136.40 0.04 0.04 0.00
Corridor Route 221.84 0.08 0.02 0.13
Louvain 1400.16 0.20 0.19 0.05
P17 SPHERE 119.90 0.01 0.00 0.02
Corridor Route 193.90 0.02 0.00 0.03
Louvain 1266.58 0.20 0.21 0.07
P18 SPHERE 123.18 0.08 0.08 0.01
Corridor Route 205.16 0.06 0.00 0.07
Louvain 1409.96 0.32 0.33 0.03
P19 SPHERE 205.97 0.02 0.02 0.00
Corridor Route 1052.15 0.40 0.30 0.31
Louvain 1401.55 0.25 0.25 0.05
P20 SPHERE 170.12 0.00 0.00 0.00
Corridor Route 391.16 0.30 0.08 0.34
Louvain 1414.17 0.10 0.08 0.04
P21 SPHERE 326.25 0.01 0.01 0.00
Corridor Route 506.65 0.20 0.14 0.12
Louvain 1502.53 0.30 0.22 0.16
P22 SPHERE 137.27 0.03 0.03 0.00
Corridor Route 180.60 0.49 0.29 0.51
Louvain 1135.72 0.28 0.26 0.08
P23 SPHERE 172.50 0.01 0.01 0.00
Corridor Route 211.28 0.08 0.04 0.11
Louvain 1065.85 0.21 0.20 0.06
P24 SPHERE 145.29 0.02 0.02 0.00
Corridor Route 202.11 0.21 0.26 0.13
Louvain 1009.77 0.27 0.28 0.07
P25 SPHERE 160.44 0.02 0.02 0.00
Corridor Route 219.89 0.11 0.04 0.14
Louvain 1059.09 0.12 0.12 0.02
P26 SPHERE 172.71 0.03 0.03 0.00
Corridor Route 203.42 0.17 0.13 0.14
Louvain 1054.63 0.24 0.19 0.13
P27 SPHERE 246.94 0.02 0.02 0.00
Corridor Route 479.93 0.14 0.06 0.11
Louvain 1388.28 0.22 0.20 0.04
P28 SPHERE 159.42 0.00 0.00 0.00
Corridor Route 267.58 0.16 0.13 0.17
Louvain 1274.86 0.19 0.16 0.08
P29 SPHERE 232.46 0.07 0.07 0.00
Corridor Route 354.39 0.14 0.11 0.07
Louvain 1219.90 0.25 0.24 0.03
P30 SPHERE 142.64 0.16 0.16 0.00
Corridor Route 175.93 0.10 0.08 0.10
Louvain 1162.54 0.15 0.09 0.10